OpenStudy (anonymous):

What is the equation of the line that passes through (2, 1) and is perpendicular to the graph of y = –x + 2?

OpenStudy (anonymous):

perpendicular is m1 x m2 = -1 and from the equation given y=-x + 2 m1 can be found to be -1 m1=gradient of y=-x+2 m2=gradient of unknown equation we want to find. -1 x m2 = -1 m2 = -1/-1 m2 = 1 use the find an equation of a line formula and we have to use the m2 found and point (2,1) y – y1 = m(x – x1) y - 1 = 1 (x-2) y - 1 = x - 2 y = x -2 + 1 y = x -1

OpenStudy (anonymous):

THANK YOU SOOO VERY MUCH

OpenStudy (anonymous):

it's ok, medal thanks :P